# Playmate: Flexible Control of Portrait Animation via 3D-Implicit Space Guided Diffusion

### v1

@@ Abstract
Recent diffusion-based talking face generation models have demonstrated impressive potential in synthesizing videos that accurately match a speech audio clip with a given reference identity. However, existing approaches still encounter significant challenges due to uncontrollable factors, such as inaccurate lip-sync, inappropriate head posture and the lack of fine-grained control over facial expressions. In order to introduce more face-guided conditions beyond speech audio clips, a novel two-stage training framework Playmate is proposed to generate more lifelike facial expressions and talking faces. In the first stage, we introduce a decoupled implicit 3D representation along with a meticulously designed motion-decoupled module to facilitate more accurate attribute disentanglement and generate expressive talking videos directly from audio cues. Then, in the second stage, we introduce an emotion-control module to encode emotion control information into the latent space, enabling fine-grained control over emotions and thereby achieving the ability to generate talking videos with desired emotion. Extensive experiments demonstrate that Playmate outperforms existing state-of-the-art methods in terms of video quality and lip-synchronization, and improves flexibility in controlling emotion and head pose. The code will be available at https://playmate.github.io.
@@
Guangzhou Quwan Network Technology
https://playmate.github.io
@@ 1 Introduction
Existing audio-driven portrait animation approaches can be broadly categorized into two main branches: GAN/NeRF-based methods (Chen et al., 2018; Prajwal et al., 2020; Lu et al., 2021; Zhang et al., 2023b; Ma et al., 2023a; Tan et al., 2024; Ki & Min, 2023; Cheng et al., 2022; Guo et al., 2021b; Wang et al., 2024; Ye et al., 2023) and diffusion-based methods (Shen et al., 2023; Sun et al., 2023; Tian et al., 2025; He et al., 2023; Xu et al., 2024b, a; Zheng et al., 2024; Cao et al., 2024; Ji et al., 2024; Ma et al., 2023b; Chen et al., 2024; Lin et al., 2024b; Liu et al., 2024a; Sun et al., 2024). The former category has primarily focused on the accuracy of lip synchronization (Chung & Zisserman, 2017) and has delivered significant results in this regard. However, GAN/NeRF-based methods often overlook the holistic coordination between audio cues and facial expressions, leading to a failure in generating expressive facial dynamics and lifelike expressions. Recently, the advent of diffusion models has facilitated the generation of high-quality images and videos (Podell et al., 2023; Esser et al., 2024; Liu et al., 2024b; Lin et al., 2024a; Kong et al., 2024; Hong et al., 2022; Yang et al., 2024). Several studies have introduced diffusion models into the field of portrait animation, enabling them to excel in generating talking face videos. Nonetheless, the expression, lip-sync, and head pose of videos generated by diffusion-based methods are strongly correlated and coupled with audio cues, which limits flexibility in controlling specific facial attributes such as head pose and emotional expression, making it challenging to modify one attribute independently without altering the associated audio content. This dependency limits the customization and adaptability of generated animations for different applications.
@@ 2.3 Facial Representation in Audio-driven Portrait Animation
Facial representation learning has been extensively studied in previous works. Various methods (Siarohin et al., 2019; Ren et al., 2021; Li et al., 2017) disentangled variables using 3DMM, sparse keypoints, or FLAME to explicitly characterize facial attributes. Furthermore, in the field of audio-driven portrait animation, several studies have introduced facial representation techniques to generate lifelike talking videos. Sadtalker (Zhang et al., 2023a) separates generation targets into different categories, including eye blinks, head poses, and lip-only 3DMM coefficients. Recent works such as VASA-1 (Xu et al., 2024b), Takin-ADA (Lin et al., 2024b), DreamTalk (Ma et al., 2023b), and JoyVASA (Cao et al., 2024) have begun to combine face representations with diffusion models to achieve more naturalistic results. Inspired by these advancements, we similarly introduce face representation techniques to generate more natural and controllable talking videos.
@@ 3.1 Expressive and Disentangled Latent Face Space Construction
Similar to VASA-1, we found that the original LivePortrait also suffered from poor disentanglement between facial dynamics and head pose. To address this issue, we introduce the pairwise head pose and facial dynamics transfer loss used in VASA-1 to improve its disentanglement. Specifically, let $I_{i}$ and $I_{j}$ be two randomly sampled frames from the same video clip. Then we transfer the head pose of $I_{j}$ onto $I_{i}$, resulting in $\hat{I}_{i,j^{pose}}=\mathcal{G}(\mathcal{W}(f_{i},x_{i},x_{i,j^{pose}}))$, the transformation is formalized as:
@@
Ultimately, a perceptual loss (Johnson et al., 2016) is employed between $\hat{I}_{i,j^{pose}}$ and $\hat{I}_{j,i^{exp}}$ o enhance the disentanglement between facial dynamics and head pose, the loss function as:
@@ 3.2 Audio Conditioned Diffusion Transformer
Speech Representation. Extensive research has demonstrated that pre-trained speech models, such as Wav2Vec2 (Baevski et al., 2020) and HuBERT (Hsu et al., 2021), outperform traditional features like MFCC in performance. Similar to other methods, we utilize Wav2Vec2 as our speech encoder to extract audio features. Additionally, we introduce a self-attention module to align audio features with motion features.
Diffusion Transformer. The architecture of the Diffusion Transformer is illustrated in Figure 2. We employ four specialized Proj modules, primarily composed of fully connected layers, to extract semantic information from four unique features, aligning this information across different modalities. Next, we utilize a multilayer Conformer (Gulati et al., 2020), a Pose-MLP, and a Exp-MLP in combination with the diffusion formulation to generate motion sequences. Diffusion models define two Markov chains: the forward chain progressively adds Gaussian noise to the target data, while the reverse chain iteratively refines the raw signal from the noise. During training, we gradually transform clean motion $m$ into Gaussian noise $m_{t}$ following the principles of Denoising Diffusion Probabilistic Models(DDPM) (Ho et al., 2020). The Diffusion Transformer is then trained to reverse this noise-adding process by taking $m_{t}$ and other conditional features as input and predicting the added noise $\epsilon$. The objective function for training can be expressed as:
@@ 4.1 Experiments Setup
Evaluation Metrics. We demonstrate the superiority of our method using multiple widely recognized metrics from previous studies. Specifically, we employ Fréchet Inception Distance (FID) (Heusel et al., 2017) and Fréchet Video Distance (FVD) (Unterthiner et al., 2019) to evaluate the quality of the generated data. For evaluating lip synchronization and motion fluidity, we compute the confidence score (Sync-C) and feature distance (Sync-D) using the pretrained SyncNet (Chung & Zisserman, 2017). Additionally, we compute the cosine similarity (CSIM) of identity vectors extracted using the ArcFace (Deng et al., 2019) face recognition model to evaluate identity preservation. Furthermore, we leverage Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018) to measure the feature-level similarity between the generated faces and their ground-truth faces.
@@ 4.2 Results and Analysis
Quantitative Results. We benchmark our method against SOTA audio-driven portrait animation methods, including Hallo (Xu et al., 2024a), Hallo2 (Cui et al., 2024), JoyVASA (Cao et al., 2024), MEMO (Zheng et al., 2024), and Sonic (Ji et al., 2024). As shown in Table 1, our Playmate significantly outperforms other methods in terms of FID, FVD, CSIM, and LPIPS on two test datasets, while also exhibiting strong competitiveness in lip synchronization. Regarding video quality, our method achieves the lowest FID and FVD scores on both test sets. Specifically, on the HDTF dataset, our FID and FVD scores are 30% and 20% lower than those of the second-best method, respectively, indicating superior video quality compared to other methods. For the CSIM and LPIPS metrics, we also achieve the best results, indicating superior performance in identity preservation and image quality. Additionally, our method achieves good results in Sync-C and Sync-D, both of which are second-best, exhibiting strong competitiveness.
Qualitative Results. Figure 4 provides a visualization comparison using open datasets. Upon analysis, previous methods were prone to generate artifacts in tooth rendering(e.g., (a)-row 6, column 3; (b)-row 3, column 1) and lip synchronization(e.g., (a)-row 4, column 7; (b)-row 2, column 7). Conversely, our approach boasts a superior decoupling capability, which allows it to create more lifelike talking head videos. For more comparison details, please see the Appendix. Since images cannot adequately reflect important aspects such as synchronization, naturalness, and stability, the full video comparison will be included in the supplementary materials or on our project URL.
@@
Emotion Control. Figure 6 shows the generation results of our method under different emotional conditions using the same audio clip. These examples clearly illustrate our model’s proficiency in encoding emotional signals into latent space and producing talking face animations with the desired emotion. This highlights Playmate’s effectiveness in achieving precise emotional control while maintaining natural and lifelike results.
@@ 4.3 Ablation Studies
CFG scales. By adjusting the CFG scales, we can strike a balance between the quality and diversity of the generated results. In Table 2, we evaluate the selection of CFG scales for the audio and emotion conditions (represented by $w_{a}$ and $w_{e}$ in Equation 8) within our model. Firstly, we conduct ablation experiments solely on $w_{a}$, as shown in the upper part of Table 2. When $w_{a}$ is set to 1.5 or 2.0, Playmate achieves better evaluation metric scores. Upon comprehensive consideration, we set $w_{a}$ to 1.5 and conduct an additional ablation experiments on $w_{e}$. In order to better measure the impact of $w_{e}$, we introduce an additional emotion classifier from (Savchenko, 2023) to calculate the emotion accuracy of the generated videos, so that we can better select an appropriate value for $w_{e}$. As shown in the bottom part of Table 2, when $w_{a}$ is set to 1.5, setting $w_{e}$ to 1.5 as well can achieve more balanced results across various testing metrics. Therefore, we also set $w_{e}$ to 1.5.
5 Conclusion
In summary, we present Playmate, a two-stage training framework designed to generate lifelike talking videos guided by speech audio clips and various optional control conditions. Our approach addresses critical limitations of existing methods, such as precise motion decoupling, expression controllability, and lip synchronization accuracy. In the first stage, Playmate utilizes a motion-decoupled module to enhance attribute disentanglement accuracy and trains a diffusion transformer to produce expressive talking videos directly from audio cues. In the second stage, an emotion-control module is introduced to encode emotion control information into the latent space, enabling fine-grained control over emotions, thereby achieving the ability to generate talking videos with desired emotion. Extensive evaluations demonstrate that Playmate consistently outperforms existing state-of-the-art solutions in terms of video quality, facial dynamics realism, and lip synchronization, while also achieving greater controllability over emotions and head pose.
Table 2: Ablation study of the audio and emotion CFG scales. The top half shows the ablation results for $w_{a}$, and the bottom half shows the ablation results when both $w_{a}$ and $w_{e}$ are combined.
@@ A.2 Gaussian Noise Augmentation
Figure 9: Ablation study of the Gaussian noise augmentation.

### v3

@@ Abstract
Recent diffusion-based talking face generation models have demonstrated impressive potential in synthesizing videos that accurately match a speech audio clip with a given reference identity. However, existing approaches still encounter significant challenges due to uncontrollable factors, such as inaccurate lip-sync, inappropriate head posture and the lack of fine-grained control over facial expressions. In order to introduce more face-guided conditions beyond speech audio clips, a novel two-stage training framework Playmate is proposed to generate more lifelike facial expressions and talking faces. In the first stage, we introduce a decoupled implicit 3D representation along with a meticulously designed motion-decoupled module to facilitate more accurate attribute disentanglement and generate expressive talking videos directly from audio cues. Then, in the second stage, we introduce an emotion-control module to encode emotion control information into the latent space, enabling fine-grained control over emotions and thereby achieving the ability to generate talking videos with desired emotion. Extensive experiments demonstrate that Playmate not only outperforms existing state-of-the-art methods in terms of video quality, but also exhibits strong competitiveness in lip synchronization while offering improved flexibility in controlling emotion and head pose. The code will be available at https://github.com/Playmate111/Playmate.
@@
Guangzhou Quwan Network Technology https://playmate.github.io/Playmate/
@@ 1 Introduction
Existing audio-driven portrait animation approaches can be broadly categorized into two main branches: GAN/NeRF-based methods (Chen et al., 2018; Prajwal et al., 2020; Lu et al., 2021; Zhang et al., 2023b; Ma et al., 2023a; Tan et al., 2024b; Ki & Min, 2023; Cheng et al., 2022; Guo et al., 2021b; Wang et al., 2024; Ye et al., 2023) and diffusion-based methods (Shen et al., 2023; Sun et al., 2023; Tian et al., 2025; He et al., 2023; Xu et al., 2024b, a; Zheng et al., 2024; Cao et al., 2024; Ji et al., 2024; Ma et al., 2023b; Chen et al., 2024; Lin et al., 2024b; Liu et al., 2024a; Sun et al., 2024). The former category has primarily focused on the accuracy of lip synchronization (Chung & Zisserman, 2017) and has delivered significant results in this regard. However, GAN/NeRF-based methods often overlook the holistic coordination between audio cues and facial expressions, leading to a failure in generating expressive facial dynamics and lifelike expressions. Recently, the advent of diffusion models has facilitated the generation of high-quality images and videos (Podell et al., 2023; Esser et al., 2024; Liu et al., 2024b; Lin et al., 2024a; Kong et al., 2024; Hong et al., 2022; Yang et al., 2024). Several studies have introduced diffusion models into the field of portrait animation, enabling them to excel in generating talking face videos. Nonetheless, the expression, lip-sync, and head pose of videos generated by diffusion-based methods are strongly correlated and coupled with audio cues, which limits flexibility in controlling specific facial attributes such as head pose and emotional expression, making it challenging to modify one attribute independently without altering the associated audio content. This dependency limits the customization and adaptability of generated animations for different applications.
@@ 2.3 Facial Representation in Audio-driven Portrait Animation
Facial representation learning has been extensively studied in previous works. Various methods (Siarohin et al., 2019; Ren et al., 2021; Li et al., 2017) disentangled variables using 3DMM, sparse keypoints, or FLAME to explicitly characterize facial attributes. Furthermore, in the field of audio-driven portrait animation, several studies have introduced facial representation techniques to generate lifelike talking videos. Sadtalker (Zhang et al., 2023a) separates generation targets into different categories, including eye blinks, head poses, and lip-only 3DMM coefficients. Recent works such as VASA-1 (Xu et al., 2024b), Takin-ADA (Lin et al., 2024b), DreamTalk (Ma et al., 2023b), and JoyVASA (Cao et al., 2024) have begun to combine face representations with diffusion models to achieve more naturalistic results. Inspired by these advancements, we similarly utilize face representation techniques to generate more natural and controllable talking videos.
@@ 3.1 Expressive and Disentangled Latent Face Space Construction
Similar to VASA-1, we found that the original LivePortrait also suffered from poor disentanglement between facial dynamics and head pose. To address this issue, we utilize the pairwise head pose and facial dynamics transfer loss used in VASA-1 to improve its disentanglement. Specifically, let $I_{i}$ and $I_{j}$ be two randomly sampled frames from the same video clip. Then we transfer the head pose of $I_{j}$ onto $I_{i}$, resulting in $\hat{I}_{i,j^{pose}}=\mathcal{G}(\mathcal{W}(f_{i},x_{i},x_{i,j^{pose}}))$, the transformation is formalized as:
@@
Ultimately, a perceptual loss (Johnson et al., 2016) is employed between $\hat{I}_{i,j^{pose}}$ and $\hat{I}_{j,i^{exp}}$ to enhance the disentanglement between facial dynamics and head pose, the loss function as:
@@ 3.2 Audio Conditioned Diffusion Transformer
Speech Representation. Extensive research has demonstrated that pre-trained speech models, such as Wav2Vec2 (Baevski et al., 2020) and HuBERT (Hsu et al., 2021), outperform traditional features like MFCC in performance. Similar to other methods, we utilize Wav2Vec2 as our speech encoder to extract audio features. Additionally, we use a self-attention module to align audio features with motion features.
Diffusion Transformer. The architecture of the Diffusion Transformer is illustrated in Figure 2. We employ four Proj modules, primarily composed of fully connected layers, to extract semantic information from four unique features, aligning this information across different modalities. Next, we utilize a multilayer Conformer (Gulati et al., 2020), a Pose-MLP, and a Exp-MLP in combination with the diffusion formulation to generate motion sequences. Diffusion models define two Markov chains: the forward chain progressively adds Gaussian noise to the target data, while the reverse chain iteratively refines the raw signal from the noise. During training, we gradually transform clean motion $m$ into Gaussian noise $m_{t}$ following the principles of Denoising Diffusion Probabilistic Models(DDPM) (Ho et al., 2020). The Diffusion Transformer is then trained to reverse this noise-adding process by taking $m_{t}$ and other conditional features as input and predicting the added noise $\epsilon$. The objective function for training can be expressed as:
@@ 4.1 Experiments Setup
Evaluation Metrics. We demonstrate the superiority of our method using multiple widely recognized metrics from previous studies. Specifically, we employ Fréchet Inception Distance (FID) (Heusel et al., 2017) and Fréchet Video Distance (FVD) (Unterthiner et al., 2019) to evaluate the quality of the generated data. For evaluating lip synchronization and motion fluidity, we compute the confidence score (Sync-C) and feature distance (Sync-D) using the pretrained SyncNet (Chung & Zisserman, 2017). Additionally, we compute the cosine similarity (CSIM) of identity vectors extracted using the ArcFace (Deng et al., 2019) face recognition model to evaluate identity preservation. Furthermore, we leverage Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018) to measure the feature-level similarity between the generated faces and their ground-truth faces. For the evaluation of emotion control, we employ an additional emotion classifier from (Savchenko, 2023) to calculate the emotion accuracy (Emo-A) of the generated videos.
@@ 4.2 Results and Analysis
Quantitative Results. We benchmark our method against SOTA audio-driven portrait animation methods, including Hallo (Xu et al., 2024a), Hallo2 (Cui et al., 2024), JoyVASA (Cao et al., 2024), MEMO (Zheng et al., 2024), and Sonic (Ji et al., 2024). As shown in Table 1, our Playmate significantly outperforms other methods in terms of FID, FVD, CSIM, and LPIPS on two test datasets, while also exhibiting strong competitiveness in lip synchronization. Regarding video quality, our method achieves the lowest FID and FVD scores on both test sets. Specifically, on the HDTF dataset, our FID and FVD scores are 30% and 20% lower than those of the second-best method, respectively, indicating superior video quality compared to other methods. For the CSIM and LPIPS metrics, we also achieve the best results, indicating superior performance in identity preservation and image quality. Additionally, our method achieves good results in Sync-C and Sync-D, both of which are second-best, exhibiting strong competitiveness. Although our quantitative metrics for lip synchronization on the two test sets are marginally lower than Sonic’s, they still outperform other compared methods. Furthermore, Sonic is a purely audio-driven algorithm that generates all features of the portrait, including lip movements, expressions, pose, etc., based on audio. This indicates that its driving flexibility is limited. In contrast, we achieve multiple controllable portrait driving methods by constructing a precise attribute disentanglement space, offering users various flexible driving options. The implementation complexity of this decoupling and subsequent driving approach is higher than that of a simple audio-driven method.
Qualitative Results. Figure 4 provides a visualization comparison using open datasets. Upon analysis, previous methods were prone to generate artifacts in tooth rendering(e.g., (a)-row 6, column 3; (b)-row 3, column 1) and lip synchronization(e.g., (a)-row 4, column 7; (b)-row 2, column 7). Conversely, our approach boasts a superior decoupling capability, which allows it to create more lifelike talking head videos. For more comparison details, please see the Appendix. Since images cannot adequately reflect important aspects such as synchronization, naturalness, and stability, the full video comparison will be included in the supplementary materials or on our project page.
@@
Table 2: Quantitative comparisons of emotion control . The best results are in bold, and the second-best are in underlined.
Methods FID $\downarrow$ FVD $\downarrow$ LPIPS $\downarrow$ Emo-A $\uparrow$
EAMM (Ji et al., 2022) 111.710 210.275 0.223 0.160
DreamTalk (Ma et al., 2023b) 119.032 199.962 0.246 0.350
EDTalk (Tan et al., 2024a) 135.215 221.897 0.289 0.460
EAT (Gan et al., 2023) 95.085 166.316 0.138 0.450
Playmate(Ours) 68.234 149.837 0.112 0.550
Emotion Control. Figure 6 shows the generation results of our method under different emotional conditions using the same audio clip. These examples clearly illustrate our model’s proficiency in encoding emotional signals into latent space and producing talking face animations with the desired emotion. This highlights Playmate’s effectiveness in achieving precise emotional control while maintaining natural and lifelike results. Furthermore, we compare our method with several existing expression control approaches across various quantitative metrics on the MEAD test dataset. As presented in Table 2, Playmate achieves the highest scores in both video quality and emotion control metrics.
Table 3: User study comparison on open dataset. The best results are in bold, and the second-best are in underlined.
Methods LS $\uparrow$ VD $\uparrow$ N $\uparrow$ VA $\uparrow$
JoyVASA (Cao et al., 2024) 2.500 2.286 1.714 1.929
Hallo (Xu et al., 2024a) 2.964 2.929 3.071 2.893
Hallo2 (Cui et al., 2024) 3.036 2.929 2.893 2.786
MEMO (Zheng et al., 2024) 3.321 3.036 3.179 3.143
Sonic (Ji et al., 2024) 3.821 3.071 3.750 3.500
Playmate(Ours) 3.750 3.857 3.464 3.643
@@
User Study. We conducted a user study involving 10 participants who rated videos using the MOS (Mean Opinion Score) rating method, on a scale of 1 to 5, across four metrics: Lip Sync (LS), Video Definition (VD), Naturalness (N), and Visual Appeal (VA). As illustrated in Table 3, Playmate has a notable advantage in the VD and VA metrics. While the LS and N metrics are slightly lower than Sonic’s, they still outperform those of other methods, showcasing Playmate’s strong competitiveness.
@@ 4.3 Ablation Studies
CFG scales. By adjusting the CFG scales, we can strike a balance between the quality and diversity of the generated results. In Table 4, we evaluate the selection of CFG scales for the audio and emotion conditions (represented by $w_{a}$ and $w_{e}$ in Equation 8) within our model. Firstly, we conduct ablation experiments solely on $w_{a}$, as shown in the upper part of Table 4. When $w_{a}$ is set to 1.5 or 2.0, Playmate achieves better evaluation metric scores. Upon comprehensive consideration, we set $w_{a}$ to 1.5 and conduct an additional ablation experiments on $w_{e}$. To more precisely assess the influence of $w_{e}$, we utilize the Emo-A metric, which helps in selecting a suitable value for $w_{e}$. As shown in the bottom part of Table 4, when $w_{a}$ is set to 1.5, setting $w_{e}$ to 1.5 as well can achieve more balanced results across various testing metrics. Therefore, we also set $w_{e}$ to 1.5.
Table 4: Ablation study of the audio and emotion CFG scales. The top half shows the ablation results for $w_{a}$, and the bottom half shows the ablation results when both $w_{a}$ and $w_{e}$ are combined. The best results are in bold, and the second-best are in underlined.
@@
5 Conclusion
In summary, we present Playmate, a two-stage training framework designed to generate lifelike talking videos guided by speech audio clips and various optional control conditions. Our approach addresses critical limitations of existing methods, such as precise motion decoupling, expression controllability, and lip synchronization accuracy. In the first stage, Playmate utilizes a motion-decoupled module to enhance attribute disentanglement accuracy and trains a diffusion transformer to produce expressive talking videos directly from audio cues. In the second stage, an emotion-control module is introduced to encode emotion control information into the latent space, enabling fine-grained control over emotions, thereby achieving the ability to generate talking videos with desired emotion. Extensive evaluations show that Playmate consistently surpasses existing state-of-the-art approaches in video quality and facial dynamics realism, while also demonstrating strong performance in lip synchronization and enhanced control over emotion and head pose.
@@ Appendix A Appendix
  3. 3.
    Evaluate the accuracy of pose conditional control. Refer to Section A.3 in the appendix.
  4. 4.
    Disentanglement of facial representation. Refer to Section A.4 in the appendix.
@@ A.2 Gaussian Noise Augmentation
Figure 9: Ablation study of the Gaussian noise augmentation.
@@
A.3 Evaluation of Pose Control
In the field of image animation, the mean keypoint distance (AKD) and the average pose distance (APD) are commonly used to evaluate the performance of pose control. We calculated the APD metric of Playmate on two datasets (HDTF and our dataset), as shown in the table below.
Table 5: Evaluation of Pose Control.
Dataset $APD_{jaw}$ $APD_{pitch}$ $APD_{roll}$
HDTF $3.003^{\circ}$ $1.308^{\circ}$ $1.214^{\circ}$
Collected dataset $3.714^{\circ}$ $1.751^{\circ}$ $1.398^{\circ}$
A.4 Disentanglement of Facial Representation
Figure 10 shows the effective disentanglement of the latents of the face. From top to bottom: the raw generated sequence, applying generated poses with fixed initial facial dynamics, and applying generated facial dynamics with fixed initial head pose and predefined spinning poses.
Figure 10: Disentanglement between head pose and facial dynamics.